OpenStudy (hba):

If cosc a =1/3t ,then tan a = ............. .

OpenStudy (dls):

cosec a=1/3t i.e sin a =3t hypotenuse=1 height=3t Find base with pythagoras. tan a = height/base.

OpenStudy (hba):

...............................

OpenStudy (dls):

?

OpenStudy (anonymous):

srry, sin a=3t cos a=rt(1-9t^2) tan a=3t/rt(1-9t^2)

OpenStudy (hba):

@Praja_01 Well you have done it right good going. :)

OpenStudy (hba):

Thanks a lot for your help @DLS

OpenStudy (dls):

np :)

OpenStudy (hba):

The answer is \[\tan a = \frac{ 3t }{ \ \sqrt{1-9t^2} }\]